Mrs Sims always claims to be five years younger than she really is. What is her real age if two and a half times her real age is three times her claimed age?

Answer by lynnlo(4176)

SIMS REAL AGE===============================X
CLAIMED AGE===================X-5
2 1/2X=3(X-5)
2 1/2X=3X-15
-1/2X=-15
X=30